A rectangular playground has a perimeter of 390 feet. The width of the playground is 75 feet. The area of the playground is how many square feet?
Answer by jim_thompson5910(35256)

We first need to find the length L

P = 2*(L + W)

390 = 2*(L + 75)

390/2 = L + 75

195 = L + 75

195-75 = L

120 = L

L = 120

The length is 120. Now use the length and width to find the area

A = L*W

A = 120*75

A = 9000

The area of the playground is 9,000 square feet